# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $0028\scriptstyle\pm 0.0049 0.0$$011\scriptstyle\pm 0.0019 0.0$$16\scriptstyle\pm 0.0028 0.000$$8\scriptstyle\pm 0.0014 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(Med) $0047\scriptstyle\pm 0.0041 0.0$$089\scriptstyle\pm 0.0069 0.0$$061\scriptstyle\pm 0.0052 0.0$$31\scriptstyle\pm 0.0026 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5145\scriptstyle\pm 0.0437$ $0.4922\scriptstyle\pm 0.0455$ $0.5028\scriptstyle\pm 0.0443$ $0.3399\scriptstyle\pm 0.0401$ $0.8444\scriptstyle\pm 0.1388$ $0.8556\scriptstyle\pm 0.0385$
L3-FT(QA) + KG $0.4871\scriptstyle\pm 0.0529$ $0.4678\scriptstyle\pm 0.0626$ $0.4771\scriptstyle\pm 0.0579$ $0.3164\scriptstyle\pm 0.0514$ $0.8667\scriptstyle\pm 0.0667$ $0.8667\scriptstyle\pm 0.0667$
G-Retriever $0.5123\scriptstyle\pm 0.0545$ $0.4100\scriptstyle\pm 0.0058$ $0.4543\scriptstyle\pm 0.0258$ $0.2958\scriptstyle\pm 0.0222$ $0.9333\scriptstyle\pm 0.0000$ $0.8778\scriptstyle\pm 0.0192$
GALAX $\mathbf{0.5385}\scriptstyle\pm\mathbf{0.0116}$ $\mathbf{0.5233}\scriptstyle\pm\mathbf{0.0200}$ $\mathbf{0.5307}\scriptstyle\pm\mathbf{0.0160}$ $\mathbf{0.3638}\scriptstyle\pm\mathbf{0.0152}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0000}$
Table 17: Model performances on ESCA
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $0028\scriptstyle\pm 0.0049 0.0$$011\scriptstyle\pm 0.0019 0.0$$16\scriptstyle\pm 0.0028 0.000$$8\scriptstyle\pm 0.0014 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(Med) $0047\scriptstyle\pm 0.0041 0.0$$089\scriptstyle\pm 0.0069 0.0$$061\scriptstyle\pm 0.0052 0.0$$31\scriptstyle\pm 0.0026 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5210\scriptstyle\pm 0.0590$ $0.4989\scriptstyle\pm 0.0691$ $0.5093\scriptstyle\pm 0.0637$ $0.3438\scriptstyle\pm 0.0577$ $0.9333\scriptstyle\pm 0.0667$ $0.9333\scriptstyle\pm 0.0000$
L3-FT(QA) + KG $0.5146\scriptstyle\pm 0.0187$ $0.4878\scriptstyle\pm 0.0051$ $0.5003\scriptstyle\pm 0.0074$ $0.3346\scriptstyle\pm 0.0059$ $0.9111\scriptstyle\pm 0.0385$ $0.9333\scriptstyle\pm 0.0000$
G-Retriever $0.5083\scriptstyle\pm 0.0142$ $0.4156\scriptstyle\pm 0.0038$ $0.4556\scriptstyle\pm 0.0076$ $0.2959\scriptstyle\pm 0.0064$ $\mathbf{0.9778}\scriptstyle\pm\mathbf{0.0385}$ $0.9444\scriptstyle\pm 0.0385$
GALAX $\mathbf{0.5775}\scriptstyle\pm\mathbf{0.0204}$ $\mathbf{0.5667}\scriptstyle\pm\mathbf{0.0260}$ $\mathbf{0.5720}\scriptstyle\pm\mathbf{0.0233}$ $\mathbf{0.4031}\scriptstyle\pm\mathbf{0.0246}$ $0.9556\scriptstyle\pm 0.0385$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0333}$
Table 18: Model performances on SCLC
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0184\scriptstyle\pm 0.0212 0.0$$389\scriptstyle\pm 0.0448 0.0$$249\scriptstyle\pm 0.0286 0.0$$131\scriptstyle\pm 0.0153 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5337\scriptstyle\pm 0.0349$ $0.4989\scriptstyle\pm 0.0479$ $0.5146\scriptstyle\pm 0.0378$ $0.3476\scriptstyle\pm 0.0341$ $0.8667\scriptstyle\pm 0.1155$ $0.8556\scriptstyle\pm 0.0694$
L3-FT(QA) + KG $0.4900\scriptstyle\pm 0.0597$ $0.4667\scriptstyle\pm 0.0677$ $0.4778\scriptstyle\pm 0.0638$ $0.3160\scriptstyle\pm 0.0555$ $0.9333\scriptstyle\pm 0.0667$ $0.8778\scriptstyle\pm 0.0694$
G-Retriever $0.4203\scriptstyle\pm 0.0263$ $0.3667\scriptstyle\pm 0.0115$ $0.3910\scriptstyle\pm 0.0180$ $0.2434\scriptstyle\pm 0.0140$ $0.8667\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.5850}\scriptstyle\pm\mathbf{0.0110}$ $\mathbf{0.5822}\scriptstyle\pm\mathbf{0.0102}$ $\mathbf{0.5836}\scriptstyle\pm\mathbf{0.0105}$ $\mathbf{0.4143}\scriptstyle\pm\mathbf{0.0102}$ $\mathbf{0.9778}\scriptstyle\pm\mathbf{0.0385}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0333}$
Table 19: Model performances on HNSC
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $0128\scriptstyle\pm 0.0222 0.0$$017\scriptstyle\pm 0.0029 0.0$$030\scriptstyle\pm 0.0051 0.0$$015\scriptstyle\pm 0.0026 0.0$$333\scriptstyle\pm 0.0577 0.0$$0167\scriptstyle\pm 0.0289$
L3-FT(Med) $0082\scriptstyle\pm 0.0036 0.0$$100\scriptstyle\pm 0.0050 0.0$$079\scriptstyle\pm 0.0016 0.0$$40\scriptstyle\pm 0.0008 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.4374\scriptstyle\pm 0.0214$ $0.4167\scriptstyle\pm 0.0208$ $0.4267\scriptstyle\pm 0.0208$ $0.2718\scriptstyle\pm 0.0174$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$
L3-FT(QA) + KG $\mathbf{0.4787}\scriptstyle\pm\mathbf{0.0445}$ $\mathbf{0.4450}\scriptstyle\pm\mathbf{0.0676}$ $\mathbf{0.4606}\scriptstyle\pm\mathbf{0.0565}$ $\mathbf{0.3006}\scriptstyle\pm\mathbf{0.0470}$ $0.9667\scriptstyle\pm 0.0577$ $0.9500\scriptstyle\pm 0.0000$
G-Retriever $0.4413\scriptstyle\pm 0.0030$ $0.3700\scriptstyle\pm 0.0000$ $0.4021\scriptstyle\pm 0.0009$ $0.2518\scriptstyle\pm 0.0008$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9500\scriptstyle\pm 0.0000$
@@
Table 20: Model performances on LUSC

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $0028\scriptstyle\pm 0.0049 0.0$$011\scriptstyle\pm 0.0019 0.0$$16\scriptstyle\pm 0.0028 0.000$$8\scriptstyle\pm 0.0014 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(Med) + Omics $0047\scriptstyle\pm 0.0041 0.0$$089\scriptstyle\pm 0.0069 0.0$$061\scriptstyle\pm 0.0052 0.0$$31\scriptstyle\pm 0.0026 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5145\scriptstyle\pm 0.0437$ $0.4922\scriptstyle\pm 0.0455$ $0.5028\scriptstyle\pm 0.0443$ $0.3399\scriptstyle\pm 0.0401$ $0.8444\scriptstyle\pm 0.1388$ $0.8556\scriptstyle\pm 0.0385$
L3-FT(CRISPR) + Omics + KG $0.4871\scriptstyle\pm 0.0529$ $0.4678\scriptstyle\pm 0.0626$ $0.4771\scriptstyle\pm 0.0579$ $0.3164\scriptstyle\pm 0.0514$ $0.8667\scriptstyle\pm 0.0667$ $0.8667\scriptstyle\pm 0.0667$
G-Retriever + pre-GAT $0.5123\scriptstyle\pm 0.0545$ $0.4100\scriptstyle\pm 0.0058$ $0.4543\scriptstyle\pm 0.0258$ $0.2958\scriptstyle\pm 0.0222$ $0.9333\scriptstyle\pm 0.0000$ $0.8778\scriptstyle\pm 0.0192$
RoG $0.5201\scriptstyle\pm 0.0255$ $0.4522\scriptstyle\pm 0.0681$ $0.4797\scriptstyle\pm 0.0520$ $0.3179\scriptstyle\pm 0.0461$ $0.8000\scriptstyle\pm 0.2309$ $0.8111\scriptstyle\pm 0.1895$
SubgraphRAG $0.4929\scriptstyle\pm 0.0245$ $0.4222\scriptstyle\pm 0.0403$ $0.4502\scriptstyle\pm 0.0304$ $0.2929\scriptstyle\pm 0.0260$ $0.6889\scriptstyle\pm 0.2143$ $0.7222\scriptstyle\pm 0.2143$
GNN-RAG $0.5298\scriptstyle\pm 0.0203$ $0.4956\scriptstyle\pm 0.0367$ $0.5092\scriptstyle\pm 0.0257$ $0.3447\scriptstyle\pm 0.0229$ $0.9556\scriptstyle\pm 0.0385$ $0.9222\scriptstyle\pm 0.0385$
GALAX $0.5385\scriptstyle\pm 0.0116$ $0.5233\scriptstyle\pm 0.0200$ $0.5307\scriptstyle\pm 0.0160$ $0.3638\scriptstyle\pm 0.0152$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0000}$
GALAX (Qwen2.5-7B) $\mathbf{0.5651}\scriptstyle\pm\mathbf{0.0206}$ $\mathbf{0.5589}\scriptstyle\pm\mathbf{0.0280}$ $\mathbf{0.5619}\scriptstyle\pm\mathbf{0.0245}$ $\mathbf{0.3938}\scriptstyle\pm\mathbf{0.0213}$ $0.8889\scriptstyle\pm 0.0770$ $0.9111\scriptstyle\pm 0.0962$
Table 25: Model performance on ESCA
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0143\scriptstyle\pm 0.0036 0.0$$178\scriptstyle\pm 0.0038 0.0$$139\scriptstyle\pm 0.0021 0.0$$71\scriptstyle\pm 0.0011 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5210\scriptstyle\pm 0.0590$ $0.4989\scriptstyle\pm 0.0691$ $0.5093\scriptstyle\pm 0.0637$ $0.3438\scriptstyle\pm 0.0577$ $0.9333\scriptstyle\pm 0.0667$ $0.9333\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics + KG $0.5146\scriptstyle\pm 0.0187$ $0.4878\scriptstyle\pm 0.0051$ $0.5003\scriptstyle\pm 0.0074$ $0.3346\scriptstyle\pm 0.0059$ $0.9111\scriptstyle\pm 0.0385$ $0.9333\scriptstyle\pm 0.0000$
G-Retriever + pre-GAT $0.5083\scriptstyle\pm 0.0142$ $0.4156\scriptstyle\pm 0.0038$ $0.4556\scriptstyle\pm 0.0076$ $0.2959\scriptstyle\pm 0.0064$ $0.9778\scriptstyle\pm 0.0385$ $0.9444\scriptstyle\pm 0.0385$
RoG $0.5275\scriptstyle\pm 0.0188$ $0.4622\scriptstyle\pm 0.0534$ $0.4877\scriptstyle\pm 0.0400$ $0.3262\scriptstyle\pm 0.0358$ $0.9111\scriptstyle\pm 0.0770$ $0.9556\scriptstyle\pm 0.0385$
SubgraphRAG $0.5373\scriptstyle\pm 0.0043$ $0.4744\scriptstyle\pm 0.0372$ $0.4990\scriptstyle\pm 0.0259$ $0.3385\scriptstyle\pm 0.0210$ $0.8889\scriptstyle\pm 0.0385$ $0.9333\scriptstyle\pm 0.0000$
GNN-RAG $0.5405\scriptstyle\pm 0.0036$ $0.5067\scriptstyle\pm 0.0404$ $0.5198\scriptstyle\pm 0.0283$ $0.3548\scriptstyle\pm 0.0243$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9778}\scriptstyle\pm\mathbf{0.0192}$
GALAX $\mathbf{0.5775}\scriptstyle\pm\mathbf{0.0204}$ $\mathbf{0.5667}\scriptstyle\pm\mathbf{0.0260}$ $\mathbf{0.5720}\scriptstyle\pm\mathbf{0.0233}$ $\mathbf{0.4031}\scriptstyle\pm\mathbf{0.0246}$ $0.9556\scriptstyle\pm 0.0385$ $0.9667\scriptstyle\pm 0.0333$
GALAX (Qwen2.5-7B) $0.5706\scriptstyle\pm 0.0228$ $0.5567\scriptstyle\pm 0.0145$ $0.5634\scriptstyle\pm 0.0173$ $0.3936\scriptstyle\pm 0.0156$ $0.9333\scriptstyle\pm 0.1155$ $0.9333\scriptstyle\pm 0.0882$
Table 26: Model performance on SCLC
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0184\scriptstyle\pm 0.0212 0.0$$389\scriptstyle\pm 0.0448 0.0$$249\scriptstyle\pm 0.0286 0.0$$131\scriptstyle\pm 0.0153 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5337\scriptstyle\pm 0.0349$ $0.4989\scriptstyle\pm 0.0479$ $0.5146\scriptstyle\pm 0.0378$ $0.3476\scriptstyle\pm 0.0341$ $0.8667\scriptstyle\pm 0.1155$ $0.8556\scriptstyle\pm 0.0694$
L3-FT(CRISPR) + Omics + KG $0.4900\scriptstyle\pm 0.0597$ $0.4667\scriptstyle\pm 0.0677$ $0.4778\scriptstyle\pm 0.0638$ $0.3160\scriptstyle\pm 0.0555$ $0.9333\scriptstyle\pm 0.0667$ $0.8778\scriptstyle\pm 0.0694$
G-Retriever + pre-GAT $0.4203\scriptstyle\pm 0.0263$ $0.3667\scriptstyle\pm 0.0115$ $0.3910\scriptstyle\pm 0.0180$ $0.2434\scriptstyle\pm 0.0140$ $0.8667\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
RoG $0.4971\scriptstyle\pm 0.1497$ $0.4689\scriptstyle\pm 0.1375$ $0.4786\scriptstyle\pm 0.1399$ $0.3313\scriptstyle\pm 0.1074$ $0.7778\scriptstyle\pm 0.2694$ $0.8000\scriptstyle\pm 0.2028$
SubgraphRAG $0.5516\scriptstyle\pm 0.0586$ $0.4578\scriptstyle\pm 0.1711$ $0.4917\scriptstyle\pm 0.1315$ $0.3346\scriptstyle\pm 0.1131$ $0.8889\scriptstyle\pm 0.1388$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.1018}$
GNN-RAG $0.5503\scriptstyle\pm 0.0732$ $0.4844\scriptstyle\pm 0.1771$ $0.5096\scriptstyle\pm 0.1386$ $0.3512\scriptstyle\pm 0.1203$ $0.8667\scriptstyle\pm 0.0000$ $0.8556\scriptstyle\pm 0.0192$
GALAX $\mathbf{0.5850}\scriptstyle\pm\mathbf{0.0110}$ $\mathbf{0.5822}\scriptstyle\pm\mathbf{0.0102}$ $\mathbf{0.5836}\scriptstyle\pm\mathbf{0.0105}$ $\mathbf{0.4143}\scriptstyle\pm\mathbf{0.0102}$ $\mathbf{0.9778}\scriptstyle\pm\mathbf{0.0385}$ $0.9000\scriptstyle\pm 0.0333$
GALAX (Qwen2.5-7B) $0.5706\scriptstyle\pm 0.0190$ $0.5667\scriptstyle\pm 0.0176$ $0.5686\scriptstyle\pm 0.0182$ $0.3988\scriptstyle\pm 0.0189$ $0.9556\scriptstyle\pm 0.0385$ $0.9000\scriptstyle\pm 0.0333$
Table 27: Model performance on HNSC
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $0128\scriptstyle\pm 0.0222 0.0$$017\scriptstyle\pm 0.0029 0.0$$029\scriptstyle\pm 0.0051 0.0$$015\scriptstyle\pm 0.0026 0.0$$333\scriptstyle\pm 0.0577 0.0$$0167\scriptstyle\pm 0.0289$
L3-FT(Med) + Omics $0082\scriptstyle\pm 0.0036 0.0$$100\scriptstyle\pm 0.0050 0.0$$079\scriptstyle\pm 0.0016 0.0$$40\scriptstyle\pm 0.0008 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.4374\scriptstyle\pm 0.0214$ $0.4167\scriptstyle\pm 0.0208$ $0.4267\scriptstyle\pm 0.0208$ $0.2718\scriptstyle\pm 0.0174$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$
L3-FT(CRISPR) + Omics + KG $0.4787\scriptstyle\pm 0.0445$ $0.4450\scriptstyle\pm 0.0676$ $0.4606\scriptstyle\pm 0.0565$ $0.3006\scriptstyle\pm 0.0470$ $0.9667\scriptstyle\pm 0.0577$ $0.9500\scriptstyle\pm 0.0000$
G-Retriever + pre-GAT $0.4413\scriptstyle\pm 0.0030$ $0.3700\scriptstyle\pm 0.0000$ $0.4021\scriptstyle\pm 0.0009$ $0.2518\scriptstyle\pm 0.0008$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9500\scriptstyle\pm 0.0000$
RoG $718\scriptstyle\pm 0.0254 0.440$$400\scriptstyle\pm 0.0673$ $0.4529\scriptstyle\pm 0.0509$ $0.2952\scriptstyle\pm 0.0404$ $0.9667\scriptstyle\pm 0.0577$ $0.9500\scriptstyle\pm 0.0500$
SubgraphRAG $0.4648\scriptstyle\pm 0.0120$ $0.4233\scriptstyle\pm 0.0635$ $0.4398\scriptstyle\pm 0.0450$ $0.2833\scriptstyle\pm 0.0355$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.0289$
GNN-RAG $0.4655\scriptstyle\pm 0.0351$ $0.4317\scriptstyle\pm 0.0813$ $0.4455\scriptstyle\pm 0.0631$ $0.2890\scriptstyle\pm 0.0506$ $0.9000\scriptstyle\pm 0.0000$ $0.9000\scriptstyle\pm 0.0000$
@@
GALAX (Qwen2.5-7B) $\mathbf{0.4787}\scriptstyle\pm\mathbf{0.0082}$ $\mathbf{0.4667}\scriptstyle\pm\mathbf{0.0153}$ $\mathbf{0.4722}\scriptstyle\pm\mathbf{0.0070}$ $\mathbf{0.3095}\scriptstyle\pm\mathbf{0.0064}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0866$
Table 28: Model performance on LUSC